OpenStudy (anonymous):

When looking at a rational function, Jamal and Angie have two different thoughts. Jamal says that the function is defined at x = −3, x = −4, and x = 6. Angie says that the function is undefined at those x values. Describe a situation where Jamal is correct, and describe a situation where Angie is correct. Is it possible for a situation to exist where they are both correct?

OpenStudy (anonymous):

For Jamal could I say he is correct if the function is (x + 3)(x + 4)(x - 6)/6

OpenStudy (anonymous):

@Directrix could you help me again?

Directrix (directrix):

From the wording in the question, do you think Jamal and Angie are both looking at the same function?

OpenStudy (anonymous):

Maybe. But it says describe a situation where they'd both be correct.

OpenStudy (anonymous):

I came up with this: Jamal would be correct if the function were (x + 3)(x + 4)(x - 6)/6 or (-3 + 3)(x + 4)(x - 6). He would be correct because the solution would be 0/6 or 0. Angie would be correct if the function were 3/(x + 3)(x + 4)(x - 6) or 3/(-3 + 3)(-4 + 4)(6 - 6). The function would be undefined at those x values because the denominator would equal 0.

Directrix (directrix):

Because if it is the same function, both J and A could not both be correct because a given function is or is not defined at specific points.

OpenStudy (anonymous):

I think I'm supposed to just come up with a function where they'd both be correct, like I did

Directrix (directrix):

y =[ (x + 3)(x + 4)(x - 6)] /6 This function is defined at x = -3, x = -4 and x = 6. How could Angie say the same function is not defined at these points?

OpenStudy (anonymous):

She could say it's not defined because the solution is 0?

Directrix (directrix):

Angie could look at the reciprocal of that function and then her statements would be true. y = 6 / [ (x + 3)(x + 4)(x - 6)]

Directrix (directrix):

0 and undefined are not the same concept.

OpenStudy (anonymous):

So I should just use the same function but flipped for Angie?

Directrix (directrix):

5/0 is undefined but 0/5 = 0 and is, of course, defined.

Directrix (directrix):

Let me think. When the function is flipped, it becomes a different function.

OpenStudy (anonymous):

I think that would make sense though. It would become a different function but I think that's fine for this question

OpenStudy (anonymous):

Because at the end it asks for a situation where they are both correct

OpenStudy (anonymous):

This is what I have now - Jamal would be correct if the function were (x + 3)(x + 4)(x - 6)/6 or (-3 + 3)(-4 + 4)(6 - 6)/6. He would be correct because the solution would be 0/6 or 0 which is defined. Angie would be correct if she looked at the reciprocal of Jamal's function which is 6/(x + 3)(x + 4)(x - 6) or 6/(-3 + 3)(-4 + 4)(6 - 6). The function would be undefined at those x values because the denominator would equal 0.

Directrix (directrix):

Look at this attachment. Let's talk about it.

OpenStudy (anonymous):

I saw that and thought it was a good example...I was trying to change mine up to avoid plagiarism

Directrix (directrix):

I don't know why people do not write y= when writing a function. Anyway, look at the function that is said to be the one where J and K are both correct.

Directrix (directrix):

I am not advocating plagiarism. I'm looking at how other people have interpreted this problem.

Directrix (directrix):

That attachment comes from here: http://openstudy.com/updates/52ad5647e4b0b09acc7f8ce5

OpenStudy (anonymous):

The function would be y = 0/0. Wouldn't this be undefined because of the denominator being 0?

Directrix (directrix):

Angie correct, Jamal wrong: y =[ (x + 3)(x + 4)(x - 6)] /6 This function is defined at x = -3, x = -4 and x = 6.

OpenStudy (anonymous):

And wouldn't Angie be incorrect there since that would be y = 0/6 which is defined?

Directrix (directrix):

This is thought of as indeterminate. 0/0 Division is defined in terms of multiplication. 6 divided by 2 = 3 if and only if 2*3 =6 5/0 is undefined because there is no number times 0 which gives 5.

Directrix (directrix):

On 0/0, any number times 0 = 0 so you could say that every number in the world could = 0/0. So, it is indeterminate. Most people just say division by 0 is undefined.

OpenStudy (anonymous):

Ah ok, so saying they are both correct if the function is y=(x + 3)(x + 4)(x - 6)/(x + 3)(x + 4)(x - 6) would be correct?

OpenStudy (anonymous):

Or not really because it's neither defined or undefined

Directrix (directrix):

Before we get to both correct, if that is true. On what equation was Jamal correct and Angie wrong?

OpenStudy (anonymous):

Jamal was correct on y=(x + 3)(x + 4)(x - 6)/6

OpenStudy (anonymous):

And Angie was wrong on that one because the function is defined

Directrix (directrix):

This function where both are supposedly correct seems off to me.

OpenStudy (anonymous):

I don't think it's possible for them to both be right on that function because it's not defined or undefined

OpenStudy (anonymous):

Was my other answer correct?

Directrix (directrix):

Many people would take y = y=[(x + 3)(x + 4)(x - 6])/[(x + 3)(x + 4)(x - 6)] and say that it is really just y = 1 after the common factors are divided out.

OpenStudy (anonymous):

That's true

Directrix (directrix):

But, that is not true because to divide, say (x+3) by (x + 3), you would have to say that the answer is y = 1 where x is not equal to -3. So, the answer is y = 1 with a hole at the point (-3, 1).

Directrix (directrix):

Y = (x^2 - 1)/(x - 1) does not have the same graph as y = x + 1. They are close but not the same.

OpenStudy (anonymous):

I'm thinking there isn't a situation where they could both be correct

Directrix (directrix):

We divide that stuff out all the time in Algebra without specifying that certain values of x have to be excluded.

Directrix (directrix):

I agree with you. I do think that there is a function where the two of them THINK they are both right. But, thinking they are right does not make their thinking correct.

Directrix (directrix):

Let's look at the problem again.

OpenStudy (anonymous):

I just don't see how a function could be defined and undefined at the same x values

Directrix (directrix):

>> Is it possible for a situation to exist where they are both correct? No, there cannot be.

OpenStudy (anonymous):

Sounds good

Directrix (directrix):

>>I just don't see how a function could be defined and undefined at the same x values Right. That is against the consistency of mathematics.

Directrix (directrix):

Whoever wrote that question is not a math person. Math people write cleanly and crisply and leave nothing to chance interpretation. So, if the computer or whatever says "wrong" on this, then we must protest.

OpenStudy (anonymous):

Well after my teacher grades it I will try to let you know what he says

OpenStudy (anonymous):

Can I show you my entire answer and we can rreview it?

Directrix (directrix):

Yes, and speak up for the mathematics we discussed and don't back down.

Directrix (directrix):

Yes, do that.

OpenStudy (anonymous):

Jamal would be correct if the function were f(x)=(x + 3)(x + 4)(x - 6)/6. He would be correct because the solution would be 0/6 or 0 which is defined. Angie would be correct if she looked at the reciprocal of Jamal's function which is f(x)=6/(x + 3)(x + 4)(x - 6). The function would be undefined at those x values because the denominator would equal 0. It is not possible for a situation to exist where they are both correct because a function can't be defined and undefined at the same x values.

Directrix (directrix):

>> He would be correct because the solution would be 0/6 or 0 which is defined. I get what you are saying but 0 is not a solution to that function. Jamal would be correct if the function were f(x)=(x + 3)(x + 4)(x - 6)/6. He would be correct because the function is defined at the points where x = - 3, -4, and 6. The value of the function at any one of those x-values would be 0.

OpenStudy (anonymous):

So just change my wording?

Directrix (directrix):

Angie would be correct if she looked at the reciprocal of Jamal's function which is f(x)=6/[(x + 3)(x + 4)(x - 6)]. The function would be undefined at those x values of -4, -4, and 6 because any one of those 3 x-values would cause the denominator to equal 0.

OpenStudy (anonymous):

ok I got it :)

Directrix (directrix):

A function is undefined at a point but a function is not just globally undefined. It is a concept of undefined at a point.

Directrix (directrix):

And, a big no on that third part. As you said, how can the function be simultaneously defined and undefined at the same points.

OpenStudy (anonymous):

Exactly. I think we've got this one figured out

Directrix (directrix):

Alrighty, then. I so enjoyed the discussion with full participation on the part of us both. Let's do it again sometime, okay?

OpenStudy (anonymous):

Okay :) Thank you again for all of your help!